### v2

# Wave–Particle Complementarity as the Optimal Limit of Unambiguous Quantum-State Discrimination

Abstract
We establish a direct operational interpretation of wave–particle complementarity in induced-coherence interferometry by linking interference visibility to the optimal performance of unambiguous quantum-state discrimination. In the low-gain Zou–Wang–Mandel interferometer, the idler modes of the two SPDC crystals act as nonorthogonal which-crystal marker states, and we show that the single-photon visibility equals the minimal inconclusive probability of the optimal Ivanovic–Dieks–Peres (IDP) strategy, $V=P_{I}^{\mathrm{opt}}$, so that the complementarity relation $D^{2}+V^{2}=1$ expresses a measurement-optimality boundary rather than a geometric constraint. Our results reveal that wave–particle duality in induced coherence is not limited by interferometer geometry but by measurement optimality in zero-error discrimination on the idler, with experimentally testable consequences predicting distinguishability beyond that inferred from visibility alone. In the presence of thermal noise we derive the hierarchy $V\leq F(\rho_{A},\rho_{B})\leq P_{I}^{\mathrm{opt}}$, linking visibility, fidelity, and optimal discrimination. These results identify distinguishability as fundamentally measurement-limited and motivate discrimination-enhanced induced-coherence imaging and sensing.
preprint: APS/123-QED
I Introduction
Interferometry based on path indistinguishability, as realized in the seminal induced-coherence experiments of Zou, Wang, and Mandel (ZWM) [wang1991induced, zou1991induced, Zou1992], demonstrates that indistinguishable idler modes erase which-way information and induce first-order coherence between the corresponding signal fields. In the low-gain regime—where at most one of the two SPDC crystals emits a photon pair per pump pulse [Shafiee2024]—single-photon interference fringes appear in the signal arm despite the idler not being detected. Beyond fundamental interest, induced coherence is crucial in quantum imaging with undetected photons [Lemos2014], induced-coherence imaging theory [Lahiri2019], and phase-sensitive sensing [Shapiro2015].
The visibility $V$ of the interference fringes and the which-way distinguishability $D$ are linked by the complementarity relation $D^{2}+V^{2}=1$ for pure two-path markers [Wootters1979, Englert1996, Jaeger1993, mohseni], and recent measurements in induced-coherence systems confirm this behavior [Torres2024]. However, in all such treatments visibility is introduced as a geometric quantity associated with the coherence term of the reduced density matrix, without a direct operational interpretation in terms of a well-defined information-processing task.
In the low-gain ZWM interferometer, each photon pair originates from either crystal A or crystal B, but never both. This defines a natural binary hypothesis test [Helstrom]: Which crystal generated the pair? The corresponding conditional idler states are nonorthogonal, so perfect discrimination is impossible. Any zero-error strategy must therefore rely on unambiguous state discrimination (USD). For two pure states, the optimal USD strategy is the Ivanovic–Dieks–Peres (IDP) three-outcome POVM [Ivanovic, Dieks, Peres, Barnett].
Distinction from prior work.— Englert’s relation $D^{2}+V^{2}\leq 1$ expresses complementarity in terms of a geometric overlap between which-path marker states [Englert1996]. For pure marker states and ideal mode overlap, as realized in the low-gain ZWM interferometer [Torres2024], this bound is saturated, $D^{2}+V^{2}=1$. Our goal is to give this equality a direct operational interpretation by identifying the conventional visibility $V$ with the optimal inconclusive probability of the Ivanovic–Dieks–Peres (IDP) unambiguous discrimination measurement, $P_{I}^{\mathrm{opt}}$, on the conditional idler states. We further show that the ZWM geometry itself realizes the optimal three-outcome IDP POVM, without ancillary modes or adaptive stages, thereby connecting complementarity to a concrete decision-theoretic task. Bagan et al. [Bagan] formulated operational duality relations using discrimination games on abstract quantum states. In contrast, we work with a specific induced-coherence interferometer, identify the physically relevant which-crystal marker states, and show that the ZWM architecture provides a direct physical implementation of the optimal IDP POVM.
[…]
Figure 1: Implementation of optimal unambiguous state discrimination (USD) [Barnett] in the induced-coherence (Zou–Wang–Mandel) interferometer [Torres2024]. Two coherently pumped nonlinear crystals (A and B) generate signal–idler photon pairs in the low-gain regime, so each emission event originates from either crystal A or crystal B but never both. The signal modes interfere at a balanced beam splitter $S_{1}$, producing single-photon fringes with visibility $V$. The idler modes ($\hat{a}_{3}^{2},\hat{a}_{3}^{3}$) form the two nonorthogonal marker states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$, whose overlap determines $V$. The output of the first idler ($\hat{a}_{3}^{2}$) beam splitter is directed to detector $D_{0}$, realizing the inconclusive IDP outcome $\Pi_{0}$. The remaining idler amplitude is processed in a Mach–Zehnder network whose internal phases and beam-splitter ratios are tuned so that detectors $D_{1}$ and $D_{2}$ project onto the orthogonal complement states $|\phi_{2}^{\perp}\rangle$ and $|\phi_{1}^{\perp}\rangle$, yielding the conclusive IDP outcomes $\Pi_{1}$ and $\Pi_{2}$. This configuration realizes the optimal IDP POVM inside the nonlinear ZWM geometry and enables an experimental test of the operational identification $P_{I}^{\mathrm{opt}}=V$ and the complementarity relation $D^{2}+V^{2}=1$.
II ZWM Idler States and Hypothesis Testing
ZWM marker states and complementarity.— In the low-gain regime [Shafiee2024, boyd, Kolobov2017], each pump pulse activates only one of the two nonlinear crystals, so each emission event corresponds to one of two mutually exclusive hypotheses. These hypotheses are encoded in the conditional idler states ($\hat{a}_{3}^{2},\hat{a}_{3}^{3}$) prepared by crystals A and B, which are generally nonorthogonal. As a result, perfect discrimination is impossible, and any zero-error identification must rely on an unambiguous state-discrimination (USD) strategy.
In the low-gain regime, the joint state generated by two coherently pumped SPDC crystals can be written schematically as
$|\Psi\rangle\approx|0\rangle+g_{1}\,|1_{S_{1}},1_{I_{1}}\rangle+g_{2}\,|1_{S_{2}},1_{I_{2}}\rangle,$ (1)
with small pair-creation amplitudes $g_{1}$ and $g_{2}$. When the idler ($\hat{a}_{3}^{2},\hat{a}_{3}^{3}$) modes are aligned to be path-identical, $I_{1}$ and $I_{2}$ become a single physical mode $I$, and the biphoton component factorizes as
$\begin{split}g_{1}|1_{S_{1}},1_{I_{1}}\rangle+g_{2}|1_{S_{2}},1_{I_{2}}\rangle\;\longrightarrow\;\\[2.0pt] |1_{I}\rangle\otimes\big(g_{1}|1_{S_{1}},0_{S_{2}}\rangle+g_{2}|0_{S_{1}},1_{S_{2}}\rangle\big),\end{split}$ (2)
so that the signal field is left in a coherent superposition even when the idler is not detected.
We now introduce an object in the shared idler $I$ path between the two crystals, modeled as a lossless beam splitter with amplitude transmittance $t$ and reflectivity $r$, $|r|^{2}+|t|^{2}=1$. In the single-pair ($\hat{a}_{1}^{1},\hat{a}_{2}^{1}$) sector, the biphoton state at the output of the two crystals can be written (see Supplemental Material) as
$\begin{split}|\psi_{si}\rangle=\frac{1}{\sqrt{2}}\Big[&|1_{S_{1}},0_{S_{2}}\rangle\big(r|1_{I_{2}},0_{I_{3}}\rangle+t|0_{I_{2}},1_{I_{3}}\rangle\big)\\ &+e^{i\phi}|0_{S_{1}},1_{S_{2}}\rangle|0_{I_{2}},1_{I_{3}}\rangle\Big],\end{split}$ (3)
where $\phi$ is the relative pump phase. Conditioned on emission in crystal A (the $|1_{S_{1}},0_{S_{2}}\rangle$ branch), the idler state is
$|\phi_{1}\rangle=r\,|1_{I_{2}},0_{I_{3}}\rangle+t\,|0_{I_{2}},1_{I_{3}}\rangle,$ (4)
while emission in crystal B (the $|0_{S_{1}},1_{S_{2}}\rangle$ branch) leaves the idler in
$|\phi_{2}\rangle=|0_{I_{2}},1_{I_{3}}\rangle.$ (5)
These idler “marker” states carry which-crystal information; their overlap is
$\langle\phi_{1}|\phi_{2}\rangle=t.$ (6)
Tracing over the idler and recombining $S_{1}$ on a balanced signal beam splitter gives single-photon interference in the signal arm with visibility
$V=|\langle\phi_{1}|\phi_{2}\rangle|=|t|.$ (7)
A natural measure of which-path distinguishability for the pure idler alternatives is
$D=\sqrt{1-|\langle\phi_{1}|\phi_{2}\rangle|^{2}}=|r|.$ (8)
Thus the ZWM interferometer realizes the standard complementarity relation
$D^{2}+V^{2}=1.$ (9)
[…]
IDP interpretation of visibility.— The two idler states in the low-gain ZWM interferometer are $|\phi_{1}\rangle=r|1,0\rangle+t|0,1\rangle$ and $|\phi_{2}\rangle=|0,1\rangle$, with overlap $|\langle\phi_{1}|\phi_{2}\rangle|=|t|=V$. In the standard complementarity framework, visibility $V$ and particle-like distinguishability
$D=\sqrt{1-V^{2}}$
are determined solely by this overlap.
We emphasize that the quantities appearing in our analysis have distinct operational meanings. Englert distinguishability $D$ is conceptually different from the two outcome probabilities of the IDP measurement. For IDP, the optimal conclusive success probability,
$P_{S}^{\mathrm{opt}}=1-|\langle\phi_{1}|\phi_{2}\rangle|,$ (10)
and the optimal inconclusive (failure) probability,
$P_{I}^{\mathrm{opt}}=|\langle\phi_{1}|\phi_{2}\rangle|,$ (11)
are complementary outcomes satisfying $P_{S}^{\mathrm{opt}}+P_{I}^{\mathrm{opt}}=1$. Thus, $P_{S}^{\mathrm{opt}}$ and $P_{I}^{\mathrm{opt}}$ are not independent quantities. Importantly, the distinguishability $D$ arises from minimum-error discrimination and is therefore not equal to either $P_{S}^{\mathrm{opt}}$ or $P_{I}^{\mathrm{opt}}$, which are the complementary outcome probabilities of unambiguous discrimination.
For pure markers in the ZWM geometry, the visibility equals the IDP failure probability,
$V=|\langle\phi_{1}|\phi_{2}\rangle|=P_{I}^{\mathrm{opt}},$ (12)
yielding the operational reformulation of complementarity,
$D^{2}+\big(P_{I}^{\mathrm{opt}}\big)^{2}=1.$ (13)
Thus, wave–particle duality in induced coherence can be understood as a statement about the optimality of a zero-error quantum measurement on the idler mode. The explicit construction of the optimal POVM $\{\Pi_{0},\Pi_{1},\Pi_{2}\}$ implementing the IDP measurement is presented in Supplemental Material.
The same reasoning applies to any balanced two-path interferometer in the single-excitation regime with pure which-path markers $|\mu_{1}\rangle$ and $|\mu_{2}\rangle$. The fringe visibility is determined by their overlap,
$V=|\langle\mu_{1}|\mu_{2}\rangle|,$
and the corresponding distinguishability is
$D=\sqrt{1-|\langle\mu_{1}|\mu_{2}\rangle|^{2}}.$
For equal priors, the optimal IDP strategy yields $P_{I}^{\mathrm{opt}}=|\langle\mu_{1}|\mu_{2}\rangle|$, so that
$D^{2}+\big(P_{I}^{\mathrm{opt}}\big)^{2}=1$
for any marker realization (polarization, spatial mode, time bin, frequency, idler photon, etc.). See supplemental material for extensions to unequal priors and mode mismatch. In this sense, complementarity for pure markers is equivalent to the fundamental limit of a zero-error state-discrimination task.
Figure 2: (a) Operational complementarity between visibility $V$, distinguishability $D$, and the optimal IDP failure probability $P_{I}^{\mathrm{opt}}$ for pure marker states with overlap $s=|\langle\mu_{1}|\mu_{2}\rangle|$ (see Appendix D). The curves satisfy $D^{2}+V^{2}=1$ and $V=P_{I}^{\mathrm{opt}}$, identifying wave–particle duality with the optimality condition of unambiguous state discrimination. (b) Comparison between distinguishability obtained via the coincidence-based method of Torres et al. [Torres2024] (solid) and via optimal IDP discrimination on the idler (dashed). The coincidence method saturates at $D_{\mathrm{coinc}}(T=1)\approx 0.52$, whereas IDP extracts greater distinguishability for intermediate $0<T<1$. Both approaches correctly give $D=0$ at $T=1$, where the marker states become identical. The advantage of IDP therefore appears only for partially overlapping markers.
Figure 2 provides a direct operational comparison between distinguishability extracted from signal-arm visibility (Torres et al.) [Torres2024] and distinguishability obtained via optimal IDP discrimination performed on the idler. In panel (a), we plot the complementarity relation for pure which-crystal markers, showing that the visibility $\mathcal{V}$, the optimal IDP failure probability $P_{I}^{\mathrm{opt}}$, and the particle-like distinguishability $D$ are constrained by the operational identity $\mathcal{V}=P_{I}^{\mathrm{opt}}$ and $D^{2}+\mathcal{V}^{2}=1$. This demonstrates that wave–particle duality is exactly equivalent to an optimal unambiguous state discrimination task for the which-path markers.
Panel (b) compares distinguishability extracted from the coincidence-based method of Torres et al. [Torres2024] (solid curve) with the distinguishability achievable via optimal Ivanovic–Dieks–Peres (IDP) unambiguous discrimination on the idler (dashed curve), both expressed in the same metric. In the Torres approach, $D$ is inferred from the imbalance between second-order correlations $g^{(2)}_{13}$ and $g^{(2)}_{23}$, and therefore saturates at a finite value $D_{\mathrm{coinc}}(T=1)<1$ even when the idler transmissivity $T$ is unity. In contrast, the IDP strategy operates directly on the which-crystal idler marker states and achieves the optimal extraction of available which-path information. Both methods converge to $D=0$ when $T=1$, since the idler marker states become identical and no measurement—IDP or otherwise—can distinguish the emission events. The advantage of IDP emerges in the intermediate regime $0<T<1$, where coincidence-based extraction underestimates the distinguishability available in principle. This highlights that path information in induced coherence is fundamentally limited by measurement strategy rather than interferometer geometry.
Operational Interpretation of Visibility as IDP Failure Probability
The result $V=P_{I}^{\mathrm{opt}}$ gives visibility a concrete information–theoretic meaning, which we now make explicit. In the induced-coherence regime, the idler modes emitted by the two SPDC crystals prepare the nonorthogonal marker states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$. Since these states satisfy $\langle\phi_{1}|\phi_{2}\rangle\neq 0$, quantum mechanics forbids any measurement from identifying the emitting crystal with zero error and nonzero success probability. The only strategy that guarantees error-free identification is unambiguous discrimination, in which the measurement is constrained to yield conclusive outcomes without ever misidentifying the state, at the cost of introducing a third, inconclusive outcome; see Appendix D. The POVM implementing this strategy has three elements $\{\Pi_{0},\Pi_{1},\Pi_{2}\}$, corresponding to conclusive identification of $|\phi_{1}\rangle$, conclusive identification of $|\phi_{2}\rangle$, and an inconclusive result.
For two pure states with equal priors, the optimal IDP measurement minimizes the inconclusive probability, which is bounded below by the state overlap itself:
$P_{I}^{\mathrm{opt}}=|\langle\phi_{1}|\phi_{2}\rangle|.$ (14)
This expression is a fundamental quantum limit: no physically allowed measurement can unambiguously discriminate $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$ with a smaller failure probability without violating the zero-error constraint. Equation (14) therefore identifies the state overlap as the operational quantity governing the difficulty of the discrimination task.
In the ZWM interferometer, the same overlap controls the visibility of single-photon interference in the signal arm, $V=|\langle\phi_{1}|\phi_{2}\rangle|$. Thus, the single-photon visibility coincides exactly with the minimum inconclusive probability of an error-free measurement on the idler:
$V=P_{I}^{\mathrm{opt}}.$ (15)
This equality shows that the conventional interference visibility is not merely a geometric measure of coherence but the optimal failure probability in a well-defined quantum decision problem. The wave–particle duality relation $D^{2}+V^{2}=1$ is therefore understood as a measurement optimality theorem: the distinguishability $D$ reflects the maximum achievable probability of conclusive identification, while the visibility $V$ corresponds to the unavoidable inconclusive probability of the optimal unambiguous discrimination protocol.
This operational perspective clarifies the role of coherence in induced interference and unifies two concepts that have historically been treated separately. Rather than simply quantifying fringe contrast, visibility characterizes the fundamental decision-theoretic limitation imposed by the nonorthogonality of the idler states. Consequently, wave–particle duality in the ZWM interferometer emerges as the boundary between error-free discrimination and coherent interference, established by the IDP optimality condition.
III Experimental Protocol
Proposed experimental implementation.— The IDP measurement can be implemented [Barnett, mohseni] directly on the idler modes of a two-crystal ZWM interferometer [Torres2024]. Two coherently pumped SPDC crystals (A,B) generate signal–idler pairs in the low-gain regime. The idler from A traverses an object $S$ with transmissivity $(t,r)$, while the idler from B bypasses it. The signal modes are combined at a balanced beam splitter to measure the single-photon visibility $V$.
To realize the IDP POVM $\{\Pi_{0},\Pi_{1},\Pi_{2}\}$ on the idler, detector $D_{1}$ implements the projection onto $|1,0\rangle$, yielding conclusive outcome $\Pi_{1}$. Detector $D_{2}$ receives the output of a variable beam splitter that interferes the idler modes with a calibrated phase such that the detected mode is $t^{*}|1,0\rangle-r^{*}|0,1\rangle$, orthogonal to $|\phi_{1}\rangle=r|1,0\rangle+t|0,1\rangle$, producing $\Pi_{2}$. The unused port defines the inconclusive outcome $\Pi_{0}$, monitored at $D_{0}$.
For each value of $|t|$, the idler counts $(N_{1},N_{2},N_{0})$ determine $P_{S}=(N_{1}+N_{2})/N_{\mathrm{tot}}$ and $P_{I}=N_{0}/N_{\mathrm{tot}}$, while the signal visibility $V$ is extracted from single-photon fringes. In the ideal regime of pure idler markers, IDP theory predicts $P_{I}=|\langle\phi_{1}|\phi_{2}\rangle|=|t|$ and $P_{S}=1-|t|$, while the ZWM analysis gives $V=|t|$ and $D=\sqrt{1-|t|^{2}}$. The experiment therefore directly tests the operational identifications $P_{I}=V$ and $D^{2}+V^{2}=1$. Details of optical layout, mode matching, loss, and detector efficiency are provided in Appendix F.
IV Mixed idler marker states and thermal noise
In realistic implementations, the idler mode that propagates between the two SPDC crystals is not perfectly isolated and may interact with environmental degrees of freedom such as thermal background radiation and loss channels [giri1, boyd]. This coupling renders the conditional idler markers associated with emission from crystals A and B mixed states, denoted by $\rho_{A}$ and $\rho_{B}$, rather than pure states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$. As a result, the interference visibility in the signal arm is no longer determined solely by the inner product of pure marker states, but is instead upper bounded by the Uhlmann fidelity $F(\rho_{A},\rho_{B})$ [Uhlmann76] between the two mixed markers; see Appendix G. By a standard purification argument one finds
$V\leq F(\rho_{A},\rho_{B}),$ (16)
which reduces to the ideal relation $V=|\langle\phi_{1}|\phi_{2}\rangle|$ in the absence of thermal noise.
From the theory of unambiguous quantum state discrimination (IDP), the optimal average failure probability $P_{I}^{\mathrm{opt}}$ for distinguishing $\rho_{A}$ from $\rho_{B}$ with equal priors satisfies
$P_{I}^{\mathrm{opt}}\geq F(\rho_{A},\rho_{B}),$ (17)
and therefore the thermal-seeded ZWM interferometer obeys the general operational duality inequality
$V\;\leq\;F(\rho_{A},\rho_{B})\;\leq\;P_{I}^{\mathrm{opt}}.$ (18)
Thus, the presence of thermal photons reduces the observed signal single-photon fringe visibility by increasing the distinguishability of the mixed conditional idler markers [giri1], while the IDP discrimination probability provides the tight operational upper bound. In the zero-noise limit these inequalities collapse to
$V=F(\rho_{A},\rho_{B})=P_{I}^{\mathrm{opt}},$ (19)
recovering the pure-state result and the standard complementarity identity $D^{2}+V^{2}=1$ in the ideal regime.
Physical intuition: The situation is closely analogous to the double-slit experiment: interference occurs only when it is impossible to know which path the photon took. In the ZWM interferometer, the two crystals play the role of the two slits, and the idler photon acts as the which-path marker. If we ignore the idler, the signal photon shows high-visibility interference. If we measure the idler strongly to learn which crystal emitted the pair, interference disappears. The key new ingredient in our proposal is the use of the Ivanovic–Dieks–Peres (IDP) unambiguous discrimination measurement, which has three possible outcomes: it sometimes identifies the emitting crystal (destroying interference) but sometimes returns an inconclusive result, in which case interference remains perfectly intact. Thus, instead of switching interference fully on or off, IDP allows us to separate the data into wave-like and particle-like subsets, revealing a direct operational meaning of visibility as the unavoidable failure probability in zero-error state discrimination. Visibility quantifies the optimal failure probability in zero-error discrimination, whereas distinguishability quantifies the optimal hypothesis-test success probability.
V Conclusion
Conclusion.— We have shown that in induced-coherence interferometry the standard complementarity relation $D^{2}+V^{2}=1$ acquires a direct operational interpretation: for pure which-crystal idler markers in the low-gain ZWM interferometer, the singles visibility equals the optimal inconclusive probability of the Ivanovic–Dieks–Peres unambiguous discrimination measurement, $V=P_{I}^{\mathrm{opt}}$. Thus wave–particle duality expresses the optimality boundary of a zero-error hypothesis test on the idler mode, demonstrating that accessible which-path information is fundamentally measurement-limited rather than interferometer-limited.
Comparison with the coincidence-based extraction of Torres et al. shows that standard correlation measurements underestimate distinguishability for partially overlapping marker states ($0<T<1$), while both approaches yield $D=0$ when the markers become identical. With thermal noise, we derived the hierarchy $V\leq F(\rho_{A},\rho_{B})\leq P_{I}^{\mathrm{opt}}$, linking visibility to Uhlmann fidelity and fidelity to optimal discrimination.
The IDP measurement naturally introduces a vacuum-flagged inconclusive outcome that functions analogously to an erasure syndrome in quantum error syndrome [Bennett1996], suggesting that control of interference can be governed by measurement design rather than physical indistinguishability [Aqua]. These results motivate discrimination-optimized induced-coherence imaging and sensing approaches [Lloyd2008, Shapiro2015, giri1], and open a route toward experiments demonstrating USD-controlled visibility.
Appendix A Schrödinger-picture derivation of the ZWM two-crystal state
In the low-gain regime of SPDC [zou1991induced], the interaction Hamiltonian is
$\hat{H}_{\mathrm{int}}=\hbar\kappa\,\hat{a}_{S}^{\dagger}\hat{a}_{I}^{\dagger}+\hbar\kappa^{*}\,\hat{a}_{S}\hat{a}_{I},$ (20)
and the corresponding weak unitary evolution generates the single-crystal state
$|\Psi_{\mathrm{single}}\rangle\approx|0\rangle+g\,|1_{S},1_{I}\rangle,$ (21)
where $g$ is a small, dimensionless gain parameter proportional to the nonlinear coupling $\kappa$ (low-gain condition $|g|\ll 1$). For two coherently pumped crystals A and B,
$\begin{split}|\psi_{SI}\rangle=&\frac{1}{\sqrt{2}}\,|1_{S_{1}},0_{S_{2}}\rangle\Big(r\,|1_{I_{2}},0_{I_{3}}\rangle+t\,|0_{I_{2}},1_{I_{3}}\rangle\Big)\\ &+\frac{1}{\sqrt{2}}\,|0_{S_{1}},1_{S_{2}}\rangle\,|0_{I_{2}},1_{I_{3}}\rangle.\end{split}$ (22)
The first term represents the creation of a signal–idler photon pair in the first nonlinear crystal, whereas the second term corresponds to pair generation in the second crystal.
A.1 Conditional idler marker states
We now define the which-crystal alternatives as [Torres2024]:
  • Crystal A emits a pair: signal state $|1_{S_{1}},0_{S_{2}}\rangle$;
  • Crystal B emits a pair: signal state $|0_{S_{1}},1_{S_{2}}\rangle$.
Conditioned on these alternatives, the corresponding idler states are:
Emission from crystal A.
Projecting the two-crystal state of Eq. (22) onto the signal branch $|1_{S_{1}},0_{S_{2}}\rangle$ isolates the contribution from crystal A. Up to normalization, the corresponding idler state is
$|\phi_{1}\rangle=r\,|1,0\rangle_{I_{2},I_{3}}+t\,|0,1\rangle_{I_{2},I_{3}}.$ (23)
where $r$ and $t$ denote the effective reflection and transmission amplitudes for the idler mode, incorporating both the pump strengths and the interferometric phase accumulation.
Emission from crystal B.
Project onto the signal branch $|0_{S_{1}},1_{S_{2}}\rangle$. The idler state is then
$|\phi_{2}\rangle=|0,1\rangle_{I_{2},I_{3}}.$ (24)
These two pure idler states encode which-crystal information. Their overlap is
$\langle\phi_{1}|\phi_{2}\rangle=r^{*}\langle 1,0|0,1\rangle+t^{*}\langle 0,1|0,1\rangle=t^{*},$ (25)
so that
$|\langle\phi_{1}|\phi_{2}\rangle|=|t|.$ (26)
A.2 Signal visibility and distinguishability
To obtain the visibility in the signal arm, we trace over the idler and recombine the signal modes $S_{1}$ on a balanced beam splitter. If both crystals are pumped [Lahirinon, Krenn2022, Shafiee2024] by the same laser with equal intensities and identical phase-matching and coupling conditions Eqs.21, the amplitudes for pair creation are $g_{1}\quad\text{and}\quad g_{2}$, and under symmetric pumping one has
$|g_{1}|=|g_{2}|.$
Conditioned on the event that a pair is generated, the a priori probabilities that the pair originated from crystal A or B are proportional to the squared amplitudes:
$\eta_{1}=\frac{|g_{1}|^{2}}{|g_{1}|^{2}+|g_{2}|^{2}},\qquad\eta_{2}=\frac{|g_{2}|^{2}}{|g_{1}|^{2}+|g_{2}|^{2}}.$ (27)
If $|g_{1}|=|g_{2}|$, then
$\eta_{1}=\eta_{2}=\frac{1}{2}.$ (28)
In particular, for equal priors $\eta_{1}=\eta_{2}=\tfrac{1}{2}$, the fringe visibility $V$ is
$V=|\langle\phi_{1}|\phi_{2}\rangle|=|t|.$ (29)
A natural measure of which-path distinguishability for two pure alternatives with equal priors is
$D=\sqrt{1-|\langle\phi_{1}|\phi_{2}\rangle|^{2}}=\sqrt{1-|t|^{2}}=|r|.$ (30)
Combining Eqs. (29) and (30), we obtain the standard complementarity relation
$D^{2}+V^{2}=(1-|t|^{2})+|t|^{2}=1.$ (31)
A.3 Imperfect mode overlap (optional extension)
If the transmitted part of the idler mode from crystal A overlaps imperfectly with the idler mode of crystal B, we introduce an overlap parameter $\gamma$ with $|\gamma|\leq 1$ such that
$|1_{I_{3}}\rangle\longrightarrow\gamma\,|1_{v_{3}},0_{w_{3}}\rangle+\sqrt{1-|\gamma|^{2}}\,|0_{v_{3}},1_{w_{3}}\rangle,$ (32)
where $v_{3}$ and $w_{3}$ are orthonormal modes. Repeating the derivation, one finds modified conditional idler states and an overlap
$\langle\phi_{1}|\phi_{2}\rangle=t\,\gamma,$ (33)
implying
$V=|\gamma|\,|t|,\qquad D=\sqrt{1-|\gamma|^{2}\,|t|^{2}}.$ (34)
The relation $D^{2}+V^{2}=1$ continues to hold in this pure-state model with imperfect mode overlap.
Appendix B Visibility and IDP failure probability for pure markers
In the low-gain ZWM configuration, the two which-crystal idler marker states are
[…]
with $|r|^{2}+|t|^{2}=1$, defined on the two-dimensional single-photon subspace of the idler output modes. Their overlap is
$|.s:=|\innerproduct{\phi_{1}}{\phi_{2}}|=|t$|. (36)
Tracing out the idler and recombining the signal modes at a balanced beam splitter yields single-photon interference fringes with visibility
[…]
and a natural distinguishability measure
$|,D=\sqrt{1-|\innerproduct{\phi_{1}}{\phi_{2}}|^{2}}=\sqrt{1-s^{2}}=|r$|, (38)
so that $D^{2}+V^{2}=1$.
[…]
The optimal Ivanovic–Dieks–Peres (IDP) strategy [Ivanovic, Dieks, Peres] minimizes the inconclusive probability
[…]
over all such POVMs. For two pure states with equal priors it is well known that the minimal value is
[…]
with the corresponding success probability $P_{S}^{\mathrm{opt}}=1-P_{I}^{\mathrm{opt}}$. For completeness, an explicit construction of the optimal POVM for the ZWM markers is given in Appendix A.
Combining Eqs. (37) and (41) we obtain
$V=P_{I}^{\mathrm{opt}},$ (42)
so that the usual complementarity relation can be written as
$D^{2}+\bigl(P_{I}^{\mathrm{opt}}\bigr)^{2}=1.$ (43)
Thus, for pure which-path markers in a balanced two-path interferometer, wave–particle duality is equivalent to the optimality condition of unambiguous quantum-state discrimination.
Appendix C Optimal IDP discrimination for the ZWM idler states
We now derive the optimal Ivanovic–Dieks–Peres (IDP) measurement [Clarke2001IDP, Danan2013Trajectories, Herzog2005MixedUSD, mohseni] for the ZWM idler states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$ defined in Eqs. (23) and (24), and we show explicitly that the optimal inconclusive probability equals the ZWM visibility.
C.1 Orthogonal complements of the idler states
The idler states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$ produced by the two crystals occupy only the single-photon idler subspace spanned by $|1,0\rangle$ and $|0,1\rangle$. In this two-dimensional subspace, each $|\phi_{j}\rangle$ admits a unique (up to phase) orthogonal complement $|\phi_{j}^{\perp}\rangle$.
We use the parametrization
$|\phi_{1}\rangle=r\,|1,0\rangle+t\,|0,1\rangle,\qquad|\phi_{2}\rangle=|0,1\rangle,$ (44)
with $|r|^{2}+|t|^{2}=1$.
The orthogonal complement of $|\phi_{2}\rangle$ is immediate:
$|\phi_{2}^{\perp}\rangle=|1,0\rangle.$ (45)
For $|\phi_{1}\rangle$, we write a general state in this subspace as $|\phi_{1}^{\perp}\rangle=\alpha\,|1,0\rangle+\beta\,|0,1\rangle$ and impose $\langle\phi_{1}^{\perp}|\phi_{1}\rangle=0$, giving
$\alpha^{*}r+\beta^{*}t=0.$ (46)
A convenient normalized solution is
$|\phi_{1}^{\perp}\rangle=t^{*}\,|1,0\rangle-r^{*}\,|0,1\rangle,$ (47)
where we used $|r|^{2}+|t|^{2}=1$.
C.2 POVM elements and zero-error conditions
The IDP measurement consists of three POVM elements $\{\Pi_{0},\Pi_{1},\Pi_{2}\}$ acting on the two-dimensional idler subspace spanned by $\{|1,0\rangle,|0,1\rangle\}$. The elements $\Pi_{1}$ and $\Pi_{2}$ correspond to conclusive identifications of $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$, respectively, while $\Pi_{0}$ corresponds to an inconclusive outcome.
Zero-error discrimination requires:
$\langle\phi_{2}|\Pi_{1}|\phi_{2}\rangle=0,\qquad\langle\phi_{1}|\Pi_{2}|\phi_{1}\rangle=0.$ (48)
These conditions are satisfied by choosing $\Pi_{1}$ and $\Pi_{2}$ to be rank-1 operators proportional to projectors onto $|\phi_{2}^{\perp}\rangle$ and $|\phi_{1}^{\perp}\rangle$:
$\Pi_{1}=\alpha\,|\phi_{2}^{\perp}\rangle\langle\phi_{2}^{\perp}|=\alpha\,|1,0\rangle\langle 1,0|,$ (49)
$\Pi_{2}=\beta\,|\phi_{1}^{\perp}\rangle\langle\phi_{1}^{\perp}|,$ (50)
with real coefficients $0\leq\alpha,\beta\leq 1$ to be determined by optimality and positivity. The inconclusive element is then
[…]
C.3 Success and inconclusive probabilities
Assuming equal a priori probabilities $\eta_{1}=\eta_{2}=\tfrac{1}{2}$ for the two idler states, the total success probability is
$P_{S}=\frac{1}{2}\,p(1|\phi_{1})+\frac{1}{2}\,p(2|\phi_{2}),$ (52)
where $p(j|\phi_{k})$ is the conditional probability of outcome $j$ given input state $|\phi_{k}\rangle$.
From Eqs. (49) and (50), we obtain
$p(1|\phi_{1})$ $=\langle\phi_{1}|\Pi_{1}|\phi_{1}\rangle=\alpha\,|\langle 1,0|\phi_{1}\rangle|^{2}=\alpha\,|r|^{2},$ (53)
$p(2|\phi_{2})$ $=\langle\phi_{2}|\Pi_{2}|\phi_{2}\rangle=\beta\,|\langle\phi_{1}^{\perp}|\phi_{2}\rangle|^{2}.$ (54)
Now
$\langle\phi_{1}^{\perp}|\phi_{2}\rangle=(t|1,0\rangle-r|0,1\rangle)\langle 0,1|=-r,$ (55)
so
$|\langle\phi_{1}^{\perp}|\phi_{2}\rangle|^{2}=|r|^{2}.$ (56)
Hence
$p(2|\phi_{2})=\beta\,|r|^{2}.$ (57)
Therefore, the average success probability is
$P_{S}=\frac{1}{2}\,\alpha|r|^{2}+\frac{1}{2}\,\beta|r|^{2}=\frac{1}{2}(\alpha+\beta)\,|r|^{2}.$ (58)
The inconclusive probability is
$P_{I}=1-P_{S}.$ (59)
C.4 Positivity of $\Pi_{0}$ and optimal coefficients
We must ensure that $\Pi_{0}$ defined by Eq. (51) is positive semidefinite. In the basis $\{|1,0\rangle,|0,1\rangle\}$, $\Pi_{1}$ is diagonal:
$\Pi_{1}=\begin{pmatrix}\alpha&0\\ 0&0\end{pmatrix}$. (60)
The operator $\Pi_{2}$ is
$\Pi_{2}=\beta\,|\phi_{1}^{\perp}\rangle\langle\phi_{1}^{\perp}|=\beta\begin{pmatrix}|t|^{2}&-tr^{*}\\ -t^{*}r&|r|^{2}\end{pmatrix}$. (61)
Hence
[…]
For $\Pi_{0}$ to be positive semidefinite, we require:
  • (i)
    The diagonal elements are nonnegative:
    $1-\alpha-\beta|t|^{2}\geq 0,\qquad 1-\beta|r|^{2}\geq 0.$ (63)
  • (ii)
    The determinant is nonnegative:
    $\det\Pi_{0}\geq 0.$ (64)
The determinant of Eq. (62) is
$\det\Pi_{0}$ $=(1-\alpha-\beta|t|^{2})(1-\beta|r|^{2})-|\beta tr^{*}|^{2}$
$=(1-\alpha-\beta|t|^{2})(1-\beta|r|^{2})-\beta^{2}|t|^{2}|r|^{2}.$ (65)
Using $|r|^{2}=1-|t|^{2}$, we set
$x:=|t|,\qquad 0\leq x\leq 1,\qquad|r|^{2}=1-x^{2}.$ (66)
To maximize $P_{S}$ in Eq. (58) subject to positivity, one finds that the optimal solution is symmetric,
$\alpha=\beta=\frac{1}{1+x}.$ (67)
Inserting $\alpha=\beta$ into $\Pi_{0}$ and demanding its smallest eigenvalue vanish (boundary of positivity) is a standard way to recover this result; the algebra is well known and we do not repeat it in full here.
Substituting $\alpha=\beta=1/(1+x)$ and $|r|^{2}=1-x^{2}$ into Eq. (58), we obtain
$P_{S}^{\mathrm{opt}}=\frac{1}{2}\cdot\frac{2}{1+x}\,(1-x^{2})=\frac{1-x^{2}}{1+x}=1-x.$ (68)
Recalling $x=|t|$ and $|\langle\phi_{1}|\phi_{2}\rangle|=|t|$, we have
$P_{S}^{\mathrm{opt}}=1-|\langle\phi_{1}|\phi_{2}\rangle|.$ (69)
Therefore, the optimal inconclusive probability is
$P_{I}^{\mathrm{opt}}=1-P_{S}^{\mathrm{opt}}=|\langle\phi_{1}|\phi_{2}\rangle|=|t|.$ (70)
Note that $P_{I}^{\mathrm{opt}}$ depends on the state overlap $|\langle\phi_{1}|\phi_{2}\rangle|$ and should not be confused with the mode occupation probability $|t|^{2}$; the IDP failure probability is set by the nonorthogonality of the states rather than by a Fock-space population.
C.5 Relation to visibility and complementarity
From Appendix A, the ZWM signal visibility is
$V=|\langle\phi_{1}|\phi_{2}\rangle|=|t|,$ (71)
and the distinguishability is
$D=\sqrt{1-|\langle\phi_{1}|\phi_{2}\rangle|^{2}}=\sqrt{1-|t|^{2}}.$ (72)
Identifying $P_{I}^{\mathrm{opt}}=|t|$, we obtain the operational identity
$D^{2}+\big(P_{I}^{\mathrm{opt}}\big)^{2}=1,$ (73)
which is the IDP interpretation of the ZWM complementarity relation used in the main text.
Appendix D Schrödinger-picture derivation of the two-crystal ZWM state
In the low-gain regime of the ZWM interferometer, alignment of the idler paths between the two crystals makes the idler modes path-identical. Denoting the two signal modes by $S_{A}$ and $S_{B}$, and the common idler mode by $I$, the single-pair component of the state can be written as
$|\Psi\rangle=\frac{1}{\sqrt{2}}\bigl(|1_{S_{A}},1_{I}\rangle+e^{i\phi}|1_{S_{B}},1_{I}\rangle\bigr),$ (74)
so that the emission alternatives “crystal A” and “crystal B” are coherently superposed but share the same idler spatial mode.
The object placed in the common idler path is modeled as a two-port beam splitter with complex transmittance $t$ and reflectance $r$, acting on the annihilation operators of the idler mode $I$ and an auxiliary vacuum mode $V$ as
$\hat{a}_{I}^{\dagger}$ $\rightarrow t\,\hat{a}_{I_{3}}^{\dagger}+r\,\hat{a}_{I_{2}}^{\dagger},$ (75)
$\hat{a}_{V}^{\dagger}$ $\rightarrow-r^{*}\,\hat{a}_{I_{3}}^{\dagger}+t^{*}\,\hat{a}_{I_{2}}^{\dagger},$ (76)
where $I_{2}$ and $I_{3}$ denote the reflected and transmitted idler modes, respectively. Acting on a single idler photon,
$|1_{I},0_{V}\rangle\;\longrightarrow\;r\,|1,0\rangle_{I_{2},I_{3}}+t\,|0,1\rangle_{I_{2},I_{3}}.$ (77)
Conditioned on emission from crystal A, the object transforms the idler into
$|\phi_{1}\rangle=r\,|1,0\rangle_{I_{2},I_{3}}+t\,|0,1\rangle_{I_{2},I_{3}},$ (78)
while emission from crystal B, which bypasses the object, prepares the idler in
$|\phi_{2}\rangle=|0,1\rangle_{I_{2},I_{3}}.$ (79)
These two pure idler states encode the which-crystal hypotheses and have overlap
$s:=\lvert\langle\phi_{1}|\phi_{2}\rangle\rvert=|t|=V,$ (80)
where $V$ is the signal singles visibility in the ZWM interferometer.
Unambiguous discrimination between $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$ with equal priors requires a three-outcome POVM $\{\Pi_{0},\Pi_{1},\Pi_{2}\}$, corresponding to conclusive identification of $|\phi_{1}\rangle$, conclusive identification of $|\phi_{2}\rangle$, and an inconclusive result. Physically, such a POVM can be realized by embedding the two-dimensional idler subspace spanned by $\{|1,0\rangle_{I_{2},I_{3}},|0,1\rangle_{I_{2},I_{3}}\}$ into a three-dimensional single-photon space by introducing a third auxiliary mode $C$ initially in the vacuum state $|0_{C}\rangle$.
In the single-photon sector of the three modes $(I_{2},I_{3},C)$ we use the basis
[…]
$|0,1,0\rangle,$
$|0,0,1\rangle\}_{I_{2},I_{3},C}.$
The Ivanovic–Dieks–Peres (IDP) measurement is implemented by a three-mode unitary $U_{\mathrm{IDP}}$ acting on the extended idler state $|\phi_{j}\rangle\otimes|0_{C}\rangle$. The most general form of $U_{\mathrm{IDP}}$ consistent with error-free discrimination is
$U_{\mathrm{IDP}}\,|\phi_{1},0_{C}\rangle$ $=\sqrt{p_{1}}\,|100\rangle+\sqrt{q_{1}}\,|001\rangle,$ (82)
$U_{\mathrm{IDP}}\,|\phi_{2},0_{C}\rangle$ $=\sqrt{p_{2}}\,|010\rangle+\sqrt{q_{2}}\,|001\rangle,$ (83)
where $p_{j}$ and $q_{j}$ ($j=1,2$) are nonnegative real numbers. In this representation
  • $p_{1}$ is the probability that input $|\phi_{1}\rangle$ produces a click in mode 1 (conclusive “crystal A”),
  • $p_{2}$ is the probability that input $|\phi_{2}\rangle$ produces a click in mode 2 (conclusive “crystal B”),
  • $q_{1}$ and $q_{2}$ are the probabilities that the respective inputs yield a click in mode 3 (inconclusive outcome).
The form of Eqs. (82)–(83) enforces the zero-error conditions: for $|\phi_{1}\rangle$ there is no amplitude in $|010\rangle$ (the “B” detector never fires), and for $|\phi_{2}\rangle$ there is no amplitude in $|100\rangle$ (the “A” detector never fires). Unitarity of $U_{\mathrm{IDP}}$ implies that the output states are normalized,
$p_{1}+q_{1}=1,\qquad p_{2}+q_{2}=1.$ (84)
Because $U_{\mathrm{IDP}}$ is unitary, it preserves inner products:
$\langle\phi_{1}|\phi_{2}\rangle=\langle\phi_{1},0_{C}|\phi_{2},0_{C}\rangle=\langle\mathrm{out}_{1}|\mathrm{out}_{2}\rangle,$ (85)
where $|\mathrm{out}_{j}\rangle=U_{\mathrm{IDP}}|\phi_{j},0_{C}\rangle$. Using Eqs. (82)–(83) and the orthogonality $\langle 100|010\rangle=\langle 100|001\rangle=\langle 010|001\rangle=0$, $\langle 001|001\rangle=1$, we find
$\langle\mathrm{out}_{1}|\mathrm{out}_{2}\rangle=\sqrt{q_{1}q_{2}},$ (86)
so that
$s=|\langle\phi_{1}|\phi_{2}\rangle|=\sqrt{q_{1}q_{2}}.$ (87)
For equal a priori probabilities and a symmetric optimal strategy, one chooses
$q_{1}=q_{2}=q,\qquad p_{1}=p_{2}=1-q,$ (88)
so that
$q=|\langle\phi_{1}|\phi_{2}\rangle|=s.$ (89)
The average success probability of the IDP measurement is then
$P_{S}^{\mathrm{opt}}=\tfrac{1}{2}p_{1}+\tfrac{1}{2}p_{2}=1-s,$ (90)
and the optimal inconclusive probability is
$P_{I}^{\mathrm{opt}}=1-P_{S}^{\mathrm{opt}}=s=|\langle\phi_{1}|\phi_{2}\rangle|.$ (91)
In the ZWM interferometer, $s=|\langle\phi_{1}|\phi_{2}\rangle|=|t|=V$, so the IDP theory yields
$P_{I}^{\mathrm{opt}}=V,\qquad P_{S}^{\mathrm{opt}}=1-V.$ (92)
Thus the conventional ZWM visibility is exactly the optimal failure probability of an error-free IDP measurement on the idler, while the which-path distinguishability is identified with the conclusive success probability. This provides the operational identity
$D^{2}+\bigl(P_{I}^{\mathrm{opt}}\bigr)^{2}=1$ (93)
for pure marker states, and establishes the interpretation $V=P_{I}^{\mathrm{opt}}$ used in the main text.
Appendix E General two-path interferometer with which-path markers
In this Appendix we provide a general derivation of the relation
$D^{2}+\big(P_{I}^{\mathrm{opt}}\big)^{2}=1$ (94)
for a balanced two-path interferometer [mohseni] in the single-excitation regime with pure which-path markers. We allow arbitrary a priori probabilities $\eta_{1},\eta_{2}$ and show how visibility, distinguishability, and optimal IDP failure probability are related.
E.1 A. Joint state and reduced density matrix
Let $S$ denote the interfering system (paths 1 and 2) and $M$ the marker system (polarization, spatial mode, idler, etc.). Conditioned on path $j\in\{1,2\}$, the marker is prepared in a pure state $|\mu_{j}\rangle$, and the probabilities of the two alternatives are $\eta_{1},\eta_{2}$ with $\eta_{1}+\eta_{2}=1$.
A convenient pure joint state is
$|\Psi\rangle=\sqrt{\eta_{1}}\,|1,0\rangle_{S}\otimes|\mu_{1}\rangle_{M}+\sqrt{\eta_{2}}\,e^{i\phi}\,|0,1\rangle_{S}\otimes|\mu_{2}\rangle_{M},$ (95)
where $\phi$ is a controllable phase. Define the marker overlap
$\gamma:=\langle\mu_{1}|\mu_{2}\rangle,\qquad|\gamma|\leq 1.$ (96)
The reduced state of the system $S$ is
$\rho_{S}=\mathrm{Tr}_{M}\left(|\Psi\rangle\langle\Psi|\right).$ (97)
Using Eq. (95), we compute
$|\Psi\rangle\langle\Psi|$ $=\eta_{1}\,|1,0\rangle\langle 1,0|\otimes|\mu_{1}\rangle\langle\mu_{1}|+\eta_{2}\,|0,1\rangle\langle 0,1|\otimes|\mu_{2}\rangle\langle\mu_{2}|$
$\quad+\sqrt{\eta_{1}\eta_{2}}e^{i\phi}|0,1\rangle\langle 1,0|\otimes|\mu_{2}\rangle\langle\mu_{1}|$
$\quad+\sqrt{\eta_{1}\eta_{2}}e^{-i\phi}|1,0\rangle\langle 0,1|\otimes|\mu_{1}\rangle\langle\mu_{2}|.$ (98)
Taking the partial trace over the meter system $M$ gives
$\mathrm{Tr}_{M}(|\mu_{1}\rangle\langle\mu_{1}|)=\mathrm{Tr}_{M}(|\mu_{2}\rangle\langle\mu_{2}|)=1,$ (99)
since $|\mu_{j}\rangle$ are normalized pure states of $M$ and their projectors have unit trace. The result is a scalar, not an operator.
$\mathrm{Tr}_{M}(|\mu_{2}\rangle\langle\mu_{1}|)=\langle\mu_{1}|\mu_{2}\rangle=\gamma.$ (100)
Thus
$\rho_{S}$ $=\eta_{1}\,|1,0\rangle\langle 1,0|+\eta_{2}\,|0,1\rangle\langle 0,1|$
$\quad+\sqrt{\eta_{1}\eta_{2}}e^{i\phi}\gamma\,|0,1\rangle\langle 1,0|+\sqrt{\eta_{1}\eta_{2}}e^{-i\phi}\gamma^{*}\,|1,0\rangle\langle 0,1|.$ (101)
In the basis $\{|1,0\rangle_{S},|0,1\rangle_{S}\}$,
$\rho_{S}=\begin{pmatrix}\eta_{1}&\sqrt{\eta_{1}\eta_{2}}e^{-i\phi}\gamma^{*}\\[4.0pt] \sqrt{\eta_{1}\eta_{2}}e^{i\phi}\gamma&\eta_{2}\end{pmatrix}$. (102)
E.2 B. Fringe visibility
For a balanced output beam splitter, the intensity at one output port varies with the phase as
$I(\varphi)=I_{0}\left[1+V\cos(\varphi+\varphi_{0})\right],$ (103)
where $V$ is the fringe visibility and $\varphi_{0}$ is a phase offset. In the single-photon regime, the visibility is related to the magnitude of the coherence (off-diagonal element) of $\rho_{S}$ as
$V=\frac{2|\rho_{01}|}{\rho_{00}+\rho_{11}}.$ (104)
Since $\rho_{00}+\rho_{11}=\eta_{1}+\eta_{2}=1$, we have
$V=2|\rho_{01}|.$ (105)
From Eq. (102),
$\rho_{01}=\sqrt{\eta_{1}\eta_{2}}e^{-i\phi}\gamma^{*},$ (106)
so
$|\rho_{01}|=\sqrt{\eta_{1}\eta_{2}}|\gamma|.$ (107)
Hence
$V=2\sqrt{\eta_{1}\eta_{2}}\,|\gamma|.$ (108)
For equal priors $\eta_{1}=\eta_{2}=\tfrac{1}{2}$, this reduces to
$V=|\gamma|=|\langle\mu_{1}|\mu_{2}\rangle|.$ (109)
E.3 C. Distinguishability
For two pure alternatives $|\mu_{1}\rangle,|\mu_{2}\rangle$ with priors $\eta_{1},\eta_{2}$, a suitable measure of which-path distinguishability is
$D=\sqrt{1-4\eta_{1}\eta_{2}|\gamma|^{2}}.$ (110)
With Eq. (108), we then have
$D^{2}+V^{2}=\big(1-4\eta_{1}\eta_{2}|\gamma|^{2}\big)+4\eta_{1}\eta_{2}|\gamma|^{2}=1.$ (111)
Thus, in this pure-state setting,
$D^{2}+V^{2}=1$ (112)
holds for arbitrary priors. For equal priors $\eta_{1}=\eta_{2}=\tfrac{1}{2}$, Eq. (110) simplifies to
$D=\sqrt{1-|\gamma|^{2}}=\sqrt{1-V^{2}}.$ (113)
E.4 D. IDP optimal failure probability
The IDP strategy for unambiguous discrimination between the two pure marker states $|\mu_{1}\rangle,|\mu_{2}\rangle$ with priors $\eta_{1},\eta_{2}$ and overlap $\gamma$ achieves an optimal inconclusive probability [Ivanovic, Dieks, Peres, Herzog2005MixedUSD]
$P_{I}^{\mathrm{opt}}=2\sqrt{\eta_{1}\eta_{2}}\,|\gamma|,$ (114)
and an optimal success probability
$P_{S}^{\mathrm{opt}}=1-P_{I}^{\mathrm{opt}}=1-2\sqrt{\eta_{1}\eta_{2}}\,|\gamma|.$ (115)
Comparing with Eq. (108), we see that
$P_{I}^{\mathrm{opt}}=V.$ (116)
Hence the distinguishability can be expressed as
$D^{2}=1-V^{2}=1-\big(P_{I}^{\mathrm{opt}}\big)^{2},$ (117)
which yields the operational relation
$D^{2}+\big(P_{I}^{\mathrm{opt}}\big)^{2}=1.$ (118)
E.5 E. Remarks on unequal priors and mixed markers
The above derivation assumes that unambiguous discrimination is possible (i.e. the supports of the two states are not nested) and that we are in the regime where the symmetric IDP solution applies. In extreme cases of highly asymmetric priors, the optimal strategy may change branch, but for the equal-pump ZWM scenario of interest ($\eta_{1}=\eta_{2}=\tfrac{1}{2}$) the expression $P_{I}^{\mathrm{opt}}=|\langle\mu_{1}|\mu_{2}\rangle|$ holds directly.
For mixed marker states $\rho_{1},\rho_{2}$, the visibility is bounded by the fidelity $F(\rho_{1},\rho_{2})$,
$V\leq F(\rho_{1},\rho_{2}),$ (119)
and the structure of optimal unambiguous discrimination becomes more involved, depending on the supports of $\rho_{1}$ and $\rho_{2}$. A full mixed-state analysis is beyond the scope of this work. In the low-gain, pure-marker regime relevant to the ZWM interferometer, Eqs. (108)–(110) and the operational identity $D^{2}+(P_{I}^{\mathrm{opt}})^{2}=1$ remain exact.
Appendix F Experimental protocol
In the IDP framework, the three-outcome measurement is described by the POVM [mohseni, Barnett] $\{\Pi_{0},\Pi_{1},\Pi_{2}\}$, where $\Pi_{1}$ and $\Pi_{2}$ correspond to conclusive identification of $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$, and $\Pi_{0}$ represents the inconclusive outcome. For a state $\rho$ entering the measurement device, the outcome probabilities follow from
$P_{i}=\mathrm{Tr}(\rho\,\Pi_{i}),\qquad i=0,1,2.$ (120)
For equal priors $\rho=\frac{1}{2}(|\phi_{1}\rangle\langle\phi_{1}|+|\phi_{2}\rangle\langle\phi_{2}|)$, the Ivanovic–Dieks–Peres theorem gives
$P_{I}^{\mathrm{opt}}=|\langle\phi_{1}|\phi_{2}\rangle|.$ (121)
In the ZWM interferometer, this overlap also sets the signal visibility, $V=|\langle\phi_{1}|\phi_{2}\rangle|$, and therefore
$P_{I}^{\mathrm{opt}}=V.$ (122)
Experimentally, the probabilities are estimated from the detected photon counts $N_{1},N_{2},N_{0}$ through relative frequencies
$P_{S}^{\mathrm{opt}}=\frac{N_{1}+N_{2}}{N_{1}+N_{2}+N_{0}},\quad P_{I}^{\mathrm{opt}}=\frac{N_{0}}{N_{1}+N_{2}+N_{0}}.$ (123)
Appendix G Mixed idler marker states with thermal noise
In the main text we worked in the ideal low-gain regime with pure idler marker states, for which the overlap $|\langle\phi_{1}|\phi_{2}\rangle|$ fixes both the single-photon visibility and the optimal IDP failure probability, $V=|\langle\phi_{1}|\phi_{2}\rangle|=P_{I}^{\mathrm{opt}}$. In this Appendix we outline how the picture is modified when thermal background photons are injected into the idler path, so that the which-crystal markers become mixed states. We work in the Schrödinger picture at the level of density operators, and make use of Uhlmann fidelity rather than repeating the full Heisenberg calculation [boyd, giri1].
G.1 A. Thermal environment and conditional idler states
We model the object $S$ in the common idler path between crystals A and B as a lossless beam splitter that couples the idler mode to a thermal background mode. Denote the corresponding annihilation operators by
$\hat{a}_{I}\;\,\text{(idler mode)},\qquad\hat{a}_{4}\;\,\text{(environment mode)}.$ (124)
The environment is taken to be in a single-mode thermal state with mean photon number $N_{B}=\langle\hat{a}_{4}^{\dagger}\hat{a}_{4}\rangle$, described by
$\rho_{\mathrm{th}}(N_{B})=\sum_{n=0}^{\infty}p_{n}\,|n\rangle\langle n|,\qquad p_{n}=\frac{N_{B}^{n}}{(1+N_{B})^{n+1}}.$ (125)
The object $S$ is represented by a beam-splitter unitary $U_{S}$ with amplitude transmittance $t_{o}$ and reflectivity $r_{o}$, $|t_{o}|^{2}+|r_{o}|^{2}=1$, acting as
$U_{S}^{\dagger}\begin{pmatrix}\hat{a}_{I_{2}}\\[2.0pt] \hat{a}_{I_{3}}\end{pmatrix}U_{S}=\begin{pmatrix}r_{o}&t_{o}\\[2.0pt] t_{o}&-r_{o}\end{pmatrix}\begin{pmatrix}\hat{a}_{I}\\[2.0pt] \hat{a}_{4}\end{pmatrix}$, (126)
where $I_{3}$ is the transmitted idler mode that seeds crystal B (and is ultimately delivered to the IDP apparatus), and $I_{2}$ is the loss port. In the single-pair regime, the relevant sector is that of at most one idler photon plus an arbitrary thermal population in mode 4.
We now define the conditional idler marker states after the object $S$, given that either crystal A or crystal B emitted the photon pair. We treat the thermal environment in the same way for both hypotheses, so that both markers are defined on the same effective output mode $I_{3}$.
Emission from crystal A.
Conditioned on emission in crystal A, the idler is in a one-photon state in mode $I$ before the object, while the thermal environment is in $\rho_{\mathrm{th}}(N_{B})$. The joint input state is therefore
$\rho^{(A)}_{\mathrm{in}}=|1_{I}\rangle\langle 1_{I}|\otimes\rho_{\mathrm{th}}(N_{B}).$ (127)
After the beam splitter $U_{S}$, the global state of modes $I_{2}$, $I_{3}$ and $4$ is
$\rho^{(A)}_{\mathrm{out}}=U_{S}\,\rho^{(A)}_{\mathrm{in}}\,U_{S}^{\dagger}.$ (128)
The idler marker associated with “crystal A fired” is defined as the reduced state of the transmitted idler mode $I_{3}$, obtained by tracing over the loss port $I_{2}$ and the remaining environment mode 4:
$\rho_{A}=\mathrm{Tr}_{I_{2},4}\big(\rho^{(A)}_{\mathrm{out}}\big)$. (129)
Because $U_{S}$ creates correlations between $I_{3}$ and the traced modes, and because the thermal state is a classical mixture, $\rho_{A}$ is in general a mixed state on the idler Hilbert space. Physically, this mixedness represents the fact that some which-path information has leaked into unobserved degrees of freedom (loss port + thermal bath).
Emission from crystal B.
Conditioned on emission in crystal B, the signal photon travels along path 2, while the idler photon is injected into the common idler path that ultimately defines mode $I_{3}$ at the IDP apparatus. To treat both hypotheses on the same footing, we also let the idler interact with the same thermal environment via $U_{S}$, but now with a possibly different effective input mode or coupling. For concreteness, we take the joint input state under hypothesis B to be
$\rho^{(B)}_{\mathrm{in}}=|\psi_{B}\rangle\langle\psi_{B}|\otimes\rho_{\mathrm{th}}(N_{B}),$ (130)
where $|\psi_{B}\rangle$ is a (generally different) single-mode pure state of the idler that reaches the object when crystal B emits. After the $U_{S}$ we have
$\rho^{(B)}_{\mathrm{out}}=U_{S}\,\rho^{(B)}_{\mathrm{in}}\,U_{S}^{\dagger},$ (131)
and the corresponding idler marker is
$\rho_{B}=\mathrm{Tr}_{I_{2},4}\big(\rho^{(B)}_{\mathrm{out}}\big)$. (132)
Again, the partial trace over the loss port and thermal mode produces a mixed state $\rho_{B}$ on mode $I_{3}$ in general. In this way, both which-crystal alternatives are represented by density operators $\rho_{A}$ and $\rho_{B}$ on the same idler output mode, and both include the effect of thermal background photons.
We emphasize that the thermal photons are not entangled with the idler in any nonclassical sense; rather, the beam splitter generates mode entanglement between $(I_{2},I_{3})$ and the environment, and tracing over the unobserved modes converts the conditional idler state into a mixed density operator.
G.2 B. Purifications, visibility, and Uhlmann fidelity
To relate the single-photon fringe visibility in the signal arm to the conditional idler markers $\rho_{A}$ and $\rho_{B}$, it is convenient to work with explicit purifications. Let $M$ denote the transmitted idler mode $I_{3}$, and let $E$ denote an abstract environment that collects the loss port $I_{2}$, the thermal mode 4, and any additional vacuum modes. By definition, each mixed idler state admits a purification on $ME$, so we introduce pure states $|\Psi_{A}\rangle_{ME}$ and $|\Psi_{B}\rangle_{ME}$ satisfying
$\rho_{A}=\mathrm{Tr}_{E}\big(|\Psi_{A}\rangle\langle\Psi_{A}|\big),\qquad\rho_{B}=\mathrm{Tr}_{E}\big(|\Psi_{B}\rangle\langle\Psi_{B}|\big).$ (133)
In the low-gain regime with equal pumping of the two crystals, the joint pure state of the signal-path qubit $S$ and the idler–environment system $ME$ is
$|\Phi\rangle_{SME}=\frac{1}{\sqrt{2}}\Big(|1\rangle_{S}\otimes|\Psi_{A}\rangle_{ME}+e^{i\phi}\,|2\rangle_{S}\otimes|\Psi_{B}\rangle_{ME}\Big),$ (134)
where $|1\rangle_{S}$ and $|2\rangle_{S}$ label the two signal paths and $\phi$ is a controllable phase.
The reduced state of the interfering signal system is obtained by tracing out the idler and environment,
$\rho_{S}=\mathrm{Tr}_{ME}\big(|\Phi\rangle\langle\Phi|\big).$ (135)
A straightforward calculation in the basis $\{|1\rangle_{S},|2\rangle_{S}\}$ gives
$\rho_{S}=\begin{pmatrix}\tfrac{1}{2}&\tfrac{1}{2}\,e^{-i\phi}\langle\Psi_{B}|\Psi_{A}\rangle\\[4.0pt] \tfrac{1}{2}\,e^{i\phi}\langle\Psi_{A}|\Psi_{B}\rangle&\tfrac{1}{2}\end{pmatrix}$. (136)
The off-diagonal element $(\rho_{S})_{12}$ determines the visibility of the single-photon interference fringes at the output of the balanced signal beam splitter. For a symmetric two-path interferometer one finds
$V=\frac{2|(\rho_{S})_{12}|}{\rho_{11}+\rho_{22}}=2\left|\frac{1}{2}\langle\Psi_{B}|\Psi_{A}\rangle\right|=|\langle\Psi_{A}|\Psi_{B}\rangle|.$ (137)
The Uhlmann fidelity between the mixed idler markers is defined as
$F(\rho_{A},\rho_{B}):=\max_{|\Psi_{A}\rangle,|\Psi_{B}\rangle}|\langle\Psi_{A}|\Psi_{B}\rangle|,$ (138)
where the maximization is over all purifications of $\rho_{A}$ and $\rho_{B}$. Comparing Eqs. (137) and (138), we immediately obtain the general bound
$V\;\leq\;F(\rho_{A},\rho_{B}),$ (139)
with equality attainable for a suitable choice of coupling between the idler and its environment. In the pure-marker limit $\rho_{A}=|\phi_{1}\rangle\langle\phi_{1}|$ and $\rho_{B}=|\phi_{2}\rangle\langle\phi_{2}|$, we can choose $|\Psi_{A}\rangle=|\phi_{1}\rangle\otimes|0_{E}\rangle$ and $|\Psi_{B}\rangle=|\phi_{2}\rangle\otimes|0_{E}\rangle$, so that $F(\rho_{A},\rho_{B})=|\langle\phi_{1}|\phi_{2}\rangle|$ and Eq. (139) reproduces the ideal relation $V=|\langle\phi_{1}|\phi_{2}\rangle|$ used in the main text.
G.3 C. Optimal IDP failure probability for mixed markers
We now connect the mixed idler markers $\rho_{A}$ and $\rho_{B}$ to optimal unambiguous state discrimination. Consider any three-outcome POVM $\{\Pi_{A},\Pi_{B},\Pi_{0}\}$ on the idler mode $I_{3}$, where $\Pi_{A}$ and $\Pi_{B}$ correspond to conclusive identification of “crystal A” and “crystal B”, respectively, and $\Pi_{0}$ represents an inconclusive outcome. For equal a priori probabilities $\eta_{A}=\eta_{B}=\tfrac{1}{2}$, the average failure probability is
[…]
subject to the zero-error constraints
[…]
The optimal IDP failure probability $P_{I}^{\mathrm{opt}}$ is obtained by minimizing $P_{I}$ over all such POVMs. For two mixed states, a general lower bound is known in terms of the Uhlmann fidelity $F(\rho_{A},\rho_{B})$ [Barnett, Herzog2005MixedUSD, Uhlmann76]:
$P_{I}^{\mathrm{opt}}\;\geq\;F(\rho_{A},\rho_{B}),$ (142)
with equality attained for pure states and for certain special classes of mixed states. In particular, for $\rho_{A}=|\phi_{1}\rangle\langle\phi_{1}|$ and $\rho_{B}=|\phi_{2}\rangle\langle\phi_{2}|$ one recovers the standard IDP result $P_{I}^{\mathrm{opt}}=|\langle\phi_{1}|\phi_{2}\rangle|$.
Combining Eqs. (139) and (142) for the thermal-seeded ZWM markers defined in Eqs. (129) and (132), we obtain the mixed-state operational duality inequality
$V\;\leq\;F(\rho_{A},\rho_{B})\;\leq\;P_{I}^{\mathrm{opt}}.$ (143)
In words: in the presence of thermal background photons, the observed signal singles visibility is always bounded above by the Uhlmann fidelity between the mixed idler markers, which in turn is bounded above by the optimal inconclusive probability of any unambiguous which-crystal measurement on the idler. In the ideal pure-marker limit (realized in the absence of thermal noise and loss), the inequalities in Eq. (143) collapse to the equality
$V=F(\rho_{A},\rho_{B})=P_{I}^{\mathrm{opt}},$ (144)
recovering the main result of the text, $V=P_{I}^{\mathrm{opt}}$, and showing that the standard complementarity relation $D^{2}+V^{2}=1$ can be written as $D^{2}+(P_{I}^{\mathrm{opt}})^{2}=1$ for the ZWM interferometer in the low-noise regime.